OpenStudy (anonymous):

Hey guys if you can help that would be great! Out of a swarm of bees one fifth part settled on a Kadamba blossom; one third on a Silihindra flower; three times the difference of those numbers flew to the bloom of a Kutaja. One bee, which remained, hovered about in the air. Tell me, charming woman, the number of bees ???????

OpenStudy (anonymous):

x/5 +x/3 +3(x/3 -x/5) +1=x 8x/15+ 3(2x/15) +1=x 14x/15=x-1 14x=15x-15 x=15

OpenStudy (anonymous):

is that the answser u have?

OpenStudy (anonymous):

ok that makes sense! Yea i tried it and it worked out the same! Thanks so much!!

OpenStudy (anonymous):

np